### v3

# Revisiting Locally Differentially Private Protocols: Towards Better Trade-offs in Privacy, Utility, and Attack Resistance

Abstract
Local Differential Privacy (LDP) offers strong privacy protection, especially in settings in which the server collecting the data is untrusted. However, designing LDP mechanisms that achieve an optimal trade-off between privacy, utility and robustness to adversarial inference and integrity attacks remains challenging. In this work, we introduce a general multi-objective optimization framework for refining LDP protocols, enabling the joint optimization of privacy and utility under various adversarial settings. While our framework is flexible to accommodate multiple privacy and security attacks as well as utility metrics, in this paper, we specifically optimize for Attacker Success Rate (ASR) under data reconstruction attack as a concrete measure of privacy leakage and Mean Squared Error (MSE) as a measure of utility. Complementarily, we evaluate integrity-oriented threats through data poisoning attacks, providing an additional adversarial perspective. More precisely, we systematically revisit these trade-offs by analyzing eight state-of-the-art LDP frequency estimation protocols and proposing refined counterparts that leverage tailored optimization techniques. Experimental results demonstrate that our proposed adaptive mechanisms consistently outperform their non-adaptive counterparts, achieving substantial reductions in ASR while preserving utility, and pushing closer to the ASR-MSE Pareto frontier. By bridging the gap between theoretical guarantees and real-world vulnerabilities, our framework enables modular and context-aware deployment of LDP mechanisms with tunable privacy-utility-attackability trade-offs.
I Introduction
Differential Privacy (DP) [1] has become a gold standard for preserving privacy in data analytics and mining, in which the goal is to ensure that an individual’s data does not significantly influence the output of any analysis. However, traditional DP relies on a trusted aggregator to apply the privacy mechanism, which is often impractical for decentralized or privacy-sensitive applications. This limitation has led to the rise of Local Differential Privacy (LDP) [2], a model that removes the need for a trusted aggregator by requiring users to perturb their data locally before sharing it with the server.
The local DP model has gained widespread adoption, with major technology companies integrating it into their systems to enhance user privacy. Notable examples include Google Chrome [3], Apple iOS [4] and Windows 10 [5], in which LDP protocols have been used to collect statistics while safeguarding individual data. A fundamental task under LDP guarantees is frequency estimation, which forms the basis of many advanced data analysis tasks like heavy hitter estimation [6, 7, 8], frequency monitoring [3, 5, 9, 10], multidimensional queries [11, 12, 13, 14], frequent item-set mining [15, 16] and spatial density estimation [17, 18].
Due to its importance, numerous LDP frequency estimation protocols have been proposed [19], namely, Generalized Randomized Response (GRR) [20], Subset Selection (SS) [21, 22], Symmetric Unary Encoding (SUE) [3], Optimized Unary Encoding (OUE) [23], Summation with Histogram Encoding (SHE) [1], Thresholding with Histogram Encoding (THE) [23], Binary Local Hashing (BLH) [6] and Optimal Local Hashing (OLH) [23]. These protocols mainly focus on improving utility, often quantified in terms of the variance (i.e., Mean Squared Error – MSE), as well as optimizing computational and communication costs [24, 25, 26] to enable efficient data collection in large-scale systems.
While utility and communication costs have been the traditional focus of LDP frequency estimation protocols, recent research has shed light on their vulnerabilities in adversarial settings. For instance, re-identification risks [27, 28] have shown that adversaries can uniquely identify users within a dataset, while attribute inference attacks [29, 30, 10] have been explored in the context of iterative data collections. Furthermore, recent works have examined data reconstruction attacks [30, 28, 31], in which adversaries aim to recover the user’s original input from its obfuscated output. Among these threats, we focus on particularly critical reconstruction attacks as they represent a worst-case leakage scenario: if an attacker can recover the user’s data from a single report, other adversaries with access to side information or repeated queries can only perform better (Examples 1 and 2).
Contributions. This paper introduces a novel attack-aware evaluation framework for analyzing and improving LDP frequency estimation protocols, designed to capture the joint trade-off between privacy leakage and utility under adversarial settings. Our approach evaluates LDP mechanisms along a unified privacy–utility–attackability trade-off curve, using the Attacker Success Rate (ASR) to quantify privacy leakage under data reconstruction attacks [30, 28, 31], and the Mean Squared Error (MSE) to measure utility [32, 23, 33]. The rationale for choosing these metrics is further discussed in Section V-B. Beyond reconstruction attacks, we further instantiate our framework to capture complementary adversarial objectives such as integrity threats through data poisoning [34], alongside ASR and MSE (see Section VI-E).
To operationalize this framework, we conduct an extensive analytical and empirical evaluation that provides new insights into the strengths and weaknesses of state-of-the-art protocols, and further enables a principled refinement of their design. In particular, we show how to tune protocol parameters through multi-objective optimization, yielding adaptive variants that better navigate the privacy-utility-attackability trade-off while preserving $\varepsilon$-LDP guarantees. Additionally, we derive the expected data reconstruction attack (i.e., analytical ASR) for three additional LDP protocols (SHE, THE and a generic Unary Encoding protocol), extending prior analyses [30, 31]. The main contributions of this paper are as follows:
  • We propose a novel evaluation framework that systematically analyzes LDP protocols under the joint lens of adversarial success (targeting privacy leakage or integrity) and utility loss. Our benchmark is designed to assist practitioners in deploying LDP protocols with customized privacy-utility-attackability trade-offs, acting as a tuning layer adaptable to various systems (e.g., telemetry [5]).
  • We derive new analytical closed-form equation of the expected data reconstruction attack for three LDP protocols, going beyond previous works [30]. These derivations enable principled reasoning about adversarial leakage and support optimization at the protocol-design level.
  • We conduct an extensive experimental evaluation of eight state-of-the-art LDP frequency estimation protocols, quantifying their behavior across a wide range of domain sizes and privacy regimes. Our results expose previously underexplored vulnerabilities and trade-offs.
  • Leveraging the insights from our analytical and empirical analyses, we propose refined adaptive variants (ASS, AUE, ALH, ATHE) of four well-known protocols (SS, OUE, OLH, THE). Our adaptive versions consistently achieve more favorable privacy-utility-attackability trade-offs, substantially reducing adversarial success rates while maintaining competitive estimation accuracy.
The code used in our experiments is available in our GitHub repository: https://github.com/hharcolezi/LDP_protocols_refined.
II Related Work
Frequency estimation is one of the primary objective of LDP that supports many downstream applications such as heavy hitter estimation [6, 7, 8], frequency monitoring [3, 5, 9, 10], multidimensional queries [11, 12, 13, 14], frequent item-set mining [15, 16], and spatial density estimation [17, 18]. A large body of work has proposed LDP mechanisms to improve estimation accuracy, reduce communication cost, or enable scalable aggregation [20, 21, 23, 22, 24, 33, 25, 26]. While these studies have established strong theoretical and practical baselines, they primarily evaluate protocols through single-objective metrics such as MSE or communication overhead.
More recently, several works have begun to examine the adversarial dimension of LDP, including data reconstruction attacks [30, 28, 31], re-identification risks [27, 28], attribute inference [29, 10, 35], and poisoning attacks [34]. These studies revealed significant vulnerabilities in existing protocols but largely stopped at qualitative observations or empirical demonstrations, without providing a unified benchmark to compare privacy leakage and utility across mechanisms.
Our work differs by taking an experimental and analytical benchmarking perspective. We quantify vulnerabilities through privacy and integrity threats and analyze their interplay with MSE across eight representative protocols. This joint evaluation highlights protocol-dependent trade-offs for privacy-utility-attackability and offers a reproducible methodology for attack-aware benchmarking. Unlike prior analyses [30], which derive ASR for a subset of mechanisms, our study extends the analytical treatment to additional protocols and demonstrates how these insights can guide the parameter tuning of existing mechanisms, yielding refined adaptive variants.
III Background
Notation. We use italic uppercase letters (e.g., $U$) to denote sets, and write $n]=\{1,\ldots,n\}$ to represent a set of $n$ positive integers. Vectors are denoted by bold lowercase letters (e.g., $\mathbf{x}$), where $\mathbf{x}_{i}$ represents the value of the $i$-th coordinate of $\mathbf{x}$. Finally, randomized mechanisms are denoted by $\mathcal{M}$}, the input domain is denoted by $\mathcal{X}$}, and the output domain by $\mathcal{Y}$}. Both $\mathcal{X}$} and $\mathcal{Y}$} are discrete, in which $|{\mathcal{X}}|=k$ and $|{\mathcal{Y}}|$ depends on the randomized mechanism $\mathcal{M}$}.
III-A Problem Statement
We consider an untrusted server collecting data from a distributed group of users while preserving their privacy. Formally, there are $n$ users, with each user holding a value $x$ from a discrete domain ${\mathcal{X}}=\{1,\ldots,k\}$. The task is frequency estimation, in which the server aims to learn the frequencies of each value across all users, denoted as $\mathbf{f}=\{f_{i}\}_{i\in[k]}$.
  • Users’ goal. Each user wants to protect their privacy. To achieve this, users apply an obfuscation mechanism $\mathcal{M}$} that perturbs their value $x$ before sending it to the server.
  • Server’s objective. The server aims to estimate the frequency distribution $\mathbf{f}$ of the values held by all users while minimizing the estimation error. After receiving the obfuscated values from all $n$ users, the server estimates a $k$-bins histogram $\hat{\mathbf{f}}=\{\hat{f}_{i}\}_{i\in[k]}$, representing the estimated frequencies.
  • Adversary’s goal. The adversary seeks to exploit the obfuscated reports to break the system guarantees, either by inferring users’ true values (i.e., data reconstruction) or by manipulating the aggregation process to bias the estimated frequencies (i.e., data poisoning).
This threat model captures a worst-case adversary (e.g., a curious or compromised server or a man-in-the-middle) who observes or injects obfuscated reports in an untrusted setting.
III-B Local Differential Privacy
Local Differential Privacy (LDP) [2] ensures that the output of a randomized mechanism does not significantly reveal information about the input. Formally:
Definition 1 ($\varepsilon$-Local Differential Privacy).
An algorithm $\mathcal{M}$} satisfies $\varepsilon$-local differential privacy ($\varepsilon$-LDP), where $\varepsilon\geq 0$, if and only if for any two distinct inputs $x,x^{\prime}\in{\mathcal{X}}$, we have:
$\forall y\in\text{Range}({\mathcal{M}}):\Pr[{\mathcal{M}}(x)=y]\leq e^{\varepsilon}\cdot\Pr[{\mathcal{M}}(x^{\prime})=y]\mathrm{,}$ (1)
in which $\text{Range}({\mathcal{M}})$ denotes the set of all possible outputs of $\mathcal{M}$}.
Smaller values of $\varepsilon$ indicate stronger privacy guarantees, as they limit how much more likely the output is for one input compared to another given the same observed value. In other words, $\varepsilon$ controls the level of indistinguishability between inputs $x$ and $x^{\prime}$, providing a formal measure of privacy.
III-C Pure LDP Framework
We consider the pure LDP framework proposed by Wang and co-authors [23] to analyze LDP frequency estimation protocols. Formally, an LDP protocol is called pure if it satisfies the following definition:
Definition 2 (Pure LDP Protocols [23]).
A protocol is considered pure if and only if there exist two probability values, $p^{*}>q^{*}$, in its perturbation mechanism $\mathcal{M}$}, such that for all inputs $x\in{\mathcal{X}}$:
$\Pr[{\mathcal{M}}(x)\in\{y\mid x\in\text{Support}(y)\}]$ $=p^{*}$
$\forall_{x^{\prime}\neq x}\Pr[{\mathcal{M}}(x^{\prime})\in\{y\mid x\in\text{Support}(y)\}]$ $=q^{*}$
in which the set $\{y\mid x\in\text{Support}(y)\}$ includes all possible outputs $y$ that “support” the input value $x$.
Thus, $p^{*}$ represents the probability that the input $x$ is mapped to an output supporting it, whereas $q^{*}$ represents the probability that any other input $x^{\prime}\neq x$ is mapped to an output that supports $x$. Given $n$ users, let $y^{j}$ denote the obfuscated value of user $j\in[n]$, the frequency estimate $\hat{f}_{i}$ of the input value $i\in[k]$ is computed as:
$\hat{f}_{i}=\frac{\sum_{j=1}^{n}\mathbb{1}_{\text{Support}(y^{j})}(i)-nq^{*}}{n(p^{*}-q^{*})}\mathrm{.}$ (2)
As shown in [23], the estimator in Eq. (2) is unbiased (i.e., $\mathbb{E}[\hat{f}_{i}]=f_{i}$) and the variance of the estimation $\hat{f}_{i}$ is:
$\mathrm{Var}[\hat{f}_{i}]=\frac{q^{*}(1-q^{*})}{n(p^{*}-q^{*})^{2}}+\frac{f_{i}(1-p^{*}-q^{*})}{n(p^{*}-q^{*})}\mathrm{.}$ (3)
With a sufficiently large domain size and no dominant frequency $f_{i}$, the second term of Eq. (3) can be ignored. Thus, as commonly used in the LDP literature [23, 9, 12], we will consider the approximate variance, which is given by:
$\mathrm{Var}[\hat{f}_{i}]=\frac{q^{*}(1-q^{*})}{n(p^{*}-q^{*})^{2}}\mathrm{.}$ (4)
Furthermore, as the estimation is unbiased, we will interchangeably refer to the variance as the MSE:
$\mathrm{MSE}=\frac{1}{k}\sum_{i=1}^{k}\mathbb{E}\left[\left(\hat{\mathbf{f}}_{i}-\mathbf{f}_{i}\right)^{2}\right]=\frac{1}{k}\sum_{i=1}^{k}\mathrm{Var}[\hat{\mathbf{f}}_{i}]\mathrm{.}$
III-D Data Reconstruction Attack on LDP
The fundamental premise of $\varepsilon$-LDP is that the input to $\mathcal{M}$} cannot be confidently inferred from its output, with the level of confidence bounded by $e^{\varepsilon}$. A user’s privacy is thus compromised if an adversary $\mathcal{A}$} can accurately infer the original input $x\in{\mathcal{X}}$ from the obfuscated output $y\in{\mathcal{Y}}$. To assess this residual inference risk, we focus on data reconstruction attacks [30, 28, 31], which directly quantify how well an adversary can recover the true input from the perturbed report.
Formally, the adversary infers:
$\hat{x}=\arg\max_{x\in{\mathcal{X}}}\Pr[x\mid y].$ (5)
Applying Bayes’ theorem gives
$\hat{x}=\arg\max_{x\in{\mathcal{X}}}\Pr[y\mid x]\cdot\Pr[x]\mathrm{.}$ (6)
Without loss of generality, we adopt a uniform prior, i.e., $\Pr[x]=1/k$ with $k=|\mathcal{X}|$, to obtain a protocol-intrinsic and distribution-agnostic measure of attackability. Thus, Eq. (6) reduces to
$\hat{x}=\arg\max_{x\in\mathcal{X}}\Pr[y\mid x],$ (7)
so that the adversary’s decision depends solely on the mechanism’s likelihood function. As shown in prior work [30], incorporating informative priors naturally increases the attack success rate by injecting side information, motivating our conservative distribution-independent choice.
We quantify reconstruction risk using the Attacker Success Rate (ASR), defined as the probability that an adversary correctly reconstructs a user’s original input [30]:
$\text{ASR}=\Pr[\hat{x}=x]=\frac{1}{n}\sum_{j=1}^{n}\mathbb{1}(\hat{x}^{j}=x^{j})\mathrm{,}$ (8)
which offers a more interpretable view of privacy protection, revealing vulnerabilities not captured by $\varepsilon$.
Connection to other inference attacks. Data reconstruction constitutes a fundamental inference task under LDP. Other attacks, such as attribute inference [29, 10, 35] and re-identification [27, 28], can be viewed as special cases that build upon reconstruction by leveraging correlations across attributes or external knowledge. This is similar to prior formulations [36, 37] in the machine learning community relating data reconstruction to attribute and membership inference.
Example 1 (Reconstruction as a proxy for re-identification).
Consider a user described by $d$ quasi-identifiers $x=(x^{(1)},\dots,x^{(d)})$, each independently perturbed under an $\varepsilon$-LDP mechanism. Let $\hat{x}^{(i)}$ denote the adversary’s reconstruction of attribute $x^{(i)}$ from the corresponding privatized report. The joint reconstruction success is defined as [28]:
$\mathrm{ASR}_{\mathrm{joint}}=\Pr[\hat{x}^{(1)}=x^{(1)}\wedge\ldots\wedge\hat{x}^{(d)}=x^{(d)}].$
Under the conditional independence of the reconstructions, this reduces to $\mathrm{ASR}_{\mathrm{joint}}=\prod_{i=1}^{d}\mathrm{ASR}_{i}$. Correctly reconstructing all quasi-identifiers might uniquely characterize a record within a population [38], making $\mathrm{ASR}_{\mathrm{joint}}$ a natural proxy for linkage attacks and re-identification.
Example 2 (Reconstruction as a proxy for attribute inference).
Consider an iterative data collection setting in which a user reports their value $x_{t}\in\mathcal{X}$ at each time step $t$ using an $\varepsilon$-LDP mechanism. Following standard deployments based on memoization (e.g., Microsoft’s ddBitFlipPM [5]), the same privatized output is reused across time as long as the underlying value remains unchanged. Let $\hat{x}_{t}$ denote the adversary’s reconstruction at time $t$. If reconstruction succeeds with non-negligible probability, the adversary can infer data changes by detecting transitions in the reconstructed outputs [10]:
$\Pr[\hat{x}_{t}\neq\hat{x}_{t-1}]\approx\Pr[x_{t}\neq x_{t-1}].$
As a result, successful reconstruction enables attribute inference over time, even when individual reports satisfy $\varepsilon$-LDP [35, 29]. This allows an adversary to infer sensitive behavioral changes (e.g., starting a diet or going on holiday) without necessarily recovering the exact value itself.
IV Overview of Pure LDP Frequency
Estimation Protocols
In this section, we provide a concise overview of eight state-of-the-art pure LDP frequency estimation protocols, selected as representative mechanisms underlying widely deployed industrial systems (e.g., SUE-based in Microsoft [5] telemetry, SS in Google Gboard [39] and OLH-based in Apple [4]). We systematically describe each mechanism based on three key functions: Encoding, Perturbation and Aggregation, which collectively define their operation. Additionally, we introduce an Attacking function for each protocol, highlighting potential vulnerabilities to privacy attacks. For protocols in which the expected ASR is not available in the literature [30], we derive it and present the detailed analyses in Appendix -A.
IV-A Generalized Randomized Response (GRR)
GRR [20] extends the randomized response method proposed by Warner [40] to a domain size of $k\geq 2$, while ensuring $\varepsilon$-LDP.
Encoding. In GRR, $\texttt{Encode}(x)=x$ and $x\in[k]$.
Perturbation. The perturbation function of GRR is:
$\Pr[\mathrm{GRR}(x)=y]=\begin{cases}p=\frac{e^{\varepsilon}}{e^{\varepsilon}+k-1}\textrm{ if }y=x,\\ q=\frac{1}{e^{\varepsilon}+k-1}\textrm{ if }y\neq x\textrm{,}\end{cases}$ (9)
in which $y\in[k]$ is the perturbed value sent to the server.
Aggregation. In GRR, each output value $i$ supports the corresponding input $i$, resulting in the support set $\mathbb{1}_{\mathrm{GRR}}={y}$. GRR is a pure protocol with $p^{*}=p$ and $q^{*}=q$. The server estimates the frequency using Eq. (2), with the following analytical MSE:
$\mathrm{MSE}_{\mathrm{GRR}}=\frac{e^{\varepsilon}+k-2}{n(e^{\varepsilon}-1)^{2}}\mathrm{.}$ (10)
Attacking. From Eq. (9), it follows that $\Pr[y=x]>\Pr[y=x^{\prime}]$ for all $x^{\prime}\in{\mathcal{X}}\setminus\{x\}$. Thus, the optimal attack strategy for GRR is to predict $\hat{x}={\mathcal{A}}_{\mathrm{GRR}}(y)=y$. The expected ASR for GRR is given by:
$\text{ASR}_{\mathrm{GRR}}=\frac{e^{\varepsilon}}{e^{\varepsilon}+k-1}\mathrm{.}$ (11)
IV-B Subset Selection (SS)
SS [21, 22] outputs a randomly selected subset $\mathbf{y}$ of size $\omega$ from the original domain $\mathcal{X}$}. SS can be seen as a generalization and optimization of GRR, in which SS is equivalent to GRR when $\omega=1$.
Encoding and Perturbation. Starting with an empty subset $\mathbf{y}=\emptyset$, the true value $x$ is added to $\mathbf{y}$ with probability: $p=\frac{\omega e^{\varepsilon}}{\omega e^{\varepsilon}+k-\omega}$. Finally, values are added to $\mathbf{y}$ as follows:
  • If $x\in\mathbf{y}$, then $\omega-1$ values are sampled from ${\mathcal{X}}\setminus\{x\}$ uniformly at random (without replacement) and are added to $\mathbf{y}$;
  • If $x\notin\mathbf{y}$, then $\omega$ values are sampled from ${\mathcal{X}}\setminus\{x\}$ uniformly at random (without replacement) and are added to $\mathbf{y}$.
The user then sends the subset $\mathbf{y}$ to the server.
Aggregation. In SS, each value $i$ in the output subset $\mathbf{y}$ supports the corresponding input value $i$. Thus, the support set for SS is $\mathbb{1}_{\mathrm{SS}}=\{x\mid x\in\mathbf{y}\}$. This protocol is pure, with: $p^{*}=p=\frac{\omega e^{\varepsilon}}{\omega e^{\varepsilon}+k-\omega}$ and $q^{*}=\frac{\omega e^{\varepsilon}(\omega-1)+(k-\omega)\omega}{(k-1)(\omega e^{\varepsilon}+k-\omega)}$. The server estimates the frequency using Eq. (2), with the following analytical MSE:
$\mathrm{MSE}_{\mathrm{SS}}=\Bigg(\frac{(k-1)(k+2\omega e^{\varepsilon}-\omega)}{n\omega\left(-k+\omega+(k-1)e^{\varepsilon}-(\omega-1)e^{\varepsilon}\right)^{2}}$ (12)
$\left.+\frac{\left(k-\omega+(\omega-1)e^{\varepsilon}\right)\left(-\omega(k-\omega)-\omega(\omega-1)e^{\varepsilon}\right)}{n\omega\left(-k+\omega+(k-1)e^{\varepsilon}-(\omega-1)e^{\varepsilon}\right)^{2}}\right)\mathrm{.}$
The optimal subset size that minimizes the MSE of SS in Eq. (12) is $\omega=\max\left(1,\left\lfloor\frac{k}{e^{\varepsilon}+1}\right\rceil\right)$ [21, 22].
Attacking. With the support set of each user’s report $\mathbb{1}_{\mathrm{SS}}$, the optimal attack strategy ${\mathcal{A}}_{\mathrm{SS}}$ is to predict $\hat{x}=\mathrm{Uniform}\left(\mathbb{1}_{\mathrm{SS}}\right)$ [30, 28]. The expected ASR for SS is [30]:
$\text{ASR}_{\mathrm{SS}}=\frac{e^{\varepsilon}}{\omega e^{\varepsilon}+k-\omega}\mathrm{.}$ (13)
IV-C Unary Encoding (UE)
UE protocols [3, 23] encode the user’s input $x\in{\mathcal{X}}$ as a $k$-dimensional one-hot vector $\mathbf{x}$, in which each bit is subsequently obfuscated independently.
Encoding. $\texttt{Encode}(x)=[0,\ldots,0,1,0,\ldots,0]$ is a binary vector with a single $1$ at position $x$ and all other positions set to 0.
Perturbation. The obfuscation function of UE mechanisms randomizes the bits from x independently to generate y as:
$\forall{i\in[k]}:\quad\Pr[\textbf{y}_{i}=1]=\begin{cases}p,\textrm{ if }\textbf{x}_{i}=1\textrm{,}\\ q,\textrm{ if }\textbf{x}_{i}=0\textrm{,}\end{cases}$ (14)
in which y is sent to the server. There are two variations of UE mechanisms: (i) Symmetric UE (SUE) [3] that selects $p=\frac{e^{\varepsilon/2}}{e^{\varepsilon/2}+1}$ and $q=\frac{1}{e^{\varepsilon/2}+1}$; and (ii) Optimized UE (OUE) [23] that selects $p=\frac{1}{2}$ and $q=\frac{1}{e^{\varepsilon}+1}$ to minimize the MSE in Eq. (15) below.
Aggregation. A reported bit vector $\mathbf{y}$ is considered to support an input $i$ if $\mathbf{y}_{i}=1$. Therefore, the support set for UE protocols is defined as $\mathbb{1}_{\mathrm{UE}}=\{i|\textbf{y}_{i}=1\}$. UE protocols are pure with $p^{*}=p$ and $q^{*}=q$. The server estimates the frequency using Eq. (2), with the corresponding MSE as:
$\mathrm{MSE}_{\mathrm{UE}}=\frac{\left(\left(e^{\varepsilon}-1\right)q+1\right)^{2}}{n\left(e^{\varepsilon}-1\right)^{2}(1-q)q}\mathrm{.}$ (15)
Attacking. Given the support set of each user’s report, $\mathbb{1}_{\mathrm{UE}}$, the adversary can adopt two possible attack strategies ${\mathcal{A}}_{\mathrm{UE}}$ [30, 28]:
  • ${\mathcal{A}}^{0}_{\mathrm{UE}}$ is a random choice $\hat{x}=\mathrm{Uniform}\left([k]\right)$, if $\mathbb{1}_{\mathrm{UE}}=\emptyset$;
  • ${\mathcal{A}}^{1}_{\mathrm{UE}}$ is a random choice $\hat{x}=\mathrm{Uniform}\left(\mathbb{1}_{\mathrm{UE}}\right)$, otherwise.
In this paper, we generalized the expected ASR of SUE and OUE given in [30] for any UE protocol as:
$\mathrm{ASR}_{\mathrm{UE}}=\frac{1}{k}\left[(1-p)(1-q)^{k-1}+\frac{p}{q}\left(1-(1-q)^{k}\right)\right]$ (16)
We defer the derivation of Eq. (16) to Appendix -A1.
IV-D Local Hashing (LH)
LH protocols [23, 6] use hash functions to map the input data $x\in{\mathcal{X}}$ to a new domain $[g]$, and then obfuscate the hash value with GRR. Let $\mathscr{H}$ be a universal hash function family such that each hash function $\mathrm{H}\in\mathscr{H}$ hashes a value $x\in{\mathcal{X}}$ into $[g]$ (i.e., $\mathrm{H}:[k]\rightarrow[g]$).
Encoding. $\texttt{Encode}(x)=\langle\mathrm{H},h\rangle$, in which $\mathrm{H}\in\mathscr{H}$ is chosen uniformly at random, and $h=\mathrm{H}(x)$. There are two variations of LH mechanisms: (i) Binary LH (BLH) [6] that just sets $g=2$, and (ii) Optimized LH (OLH) [23] that selects $g=\lfloor e^{\varepsilon}+1\rceil$ to minimize the MSE in Eq. (18) below.
Perturbation. LH protocols perturb $\langle\mathrm{H},h\rangle$ into $\langle\mathrm{H},y\rangle$, just like GRR, as follows:
$\forall i\in[g],\;\Pr[y=i]=\begin{cases}p=\frac{e^{\varepsilon}}{e^{\varepsilon}+g-1},&\text{if }h=i\mathrm{,}\\ q=\frac{1}{e^{\varepsilon}+g-1},&\text{if }h\neq i\mathrm{.}\end{cases}$ (17)
Aggregation. For each reported tuple $\langle\mathrm{H},y\rangle$, the support set for LH protocols consists of all values $x\in{\mathcal{X}}$ that hash to $y$, denoted as $\mathbb{1}_{\mathrm{LH}}=\{x\mid\mathrm{H}(x)=y\}$. LH protocols are pure with $p^{*}=p$ and $q^{*}=\frac{1}{g}$. The server estimates the frequency using Eq. (2) with the following analytical MSE:
$\mathrm{MSE}_{\mathrm{LH}}=\frac{(e^{\varepsilon}-1+g)^{2}}{n(e^{\varepsilon}-1)^{2}(g-1)}\mathrm{.}$ (18)
Attacking. Based on the support set of each user’s report, $\mathbb{1}_{\mathrm{LH}}$, the adversary can employ one of two possible attack strategies, denoted by ${\mathcal{A}}_{\mathrm{LH}}$ [30, 28]:
  • ${\mathcal{A}}^{0}_{\mathrm{LH}}$ is a random choice $\hat{x}=\mathrm{Uniform}\left([k]\right)$, if $\mathbb{1}_{\mathrm{LH}}=\emptyset$;
  • ${\mathcal{A}}^{1}_{\mathrm{LH}}$ is a random choice $\hat{x}=\mathrm{Uniform}\left(\mathbb{1}_{\mathrm{LH}}\right)$, otherwise.
The expected ASR of LH protocols is given by [30]:
$\text{ASR}_{\mathrm{LH}}=\frac{e^{\varepsilon}}{(e^{\varepsilon}+g-1)\cdot\max\left\{\frac{k}{g},1\right\}}\mathrm{.}$ (19)
IV-E Histogram Encoding (HE)
HE protocols [23] encode the user’s input data $x\in{\mathcal{X}}$, as a one-hot $k$-dimensional histogram before obfuscating each bit independently.
Encoding. $\texttt{Encode}(x)=[0.0,0.0,\ldots,1.0,0.0,\ldots,0.0]$ in which only the $x$-th component is $1.0$. Two different input values $x,x^{\prime}\in{\mathcal{X}}$ will result in two vectors with L1 distance of $\Delta_{1}=2$.
Perturbation. The perturbation function $\texttt{Perturb}(\mathbf{x})$ generates the output vector $\mathbf{y}$, where each component is given by $\textbf{y}_{i}=\textbf{x}_{i}+\mathrm{Lap}\left(\frac{\Delta_{1}}{\varepsilon}\right)$, with $\mathrm{Lap}(\cdot)$ representing the Laplace mechanism [1].
The following subsections describe two HE-based mechanisms: Summation with HE (SHE) and Thresholding with HE (THE). These mechanisms differ in their aggregation and attack strategies.
IV-E1 Summation with HE (SHE)
With SHE, there is no post-processing of y at the server side.
Aggregation. Since Laplace noise with mean 0 is added to each vector independently, the server estimates the frequency using the sum of the noisy reports: $\hat{\mathbf{f}}=\left\{\sum_{j=1}^{n}\mathbf{y}^{j}_{i}\right\}_{i\in[k]}$. This aggregation method for SHE does not provide a support set and is not pure. The analytical MSE of this estimation is:
$\mathrm{MSE}_{\mathrm{SHE}}=\frac{8}{n\varepsilon^{2}}\mathrm{.}$ (20)
Attacking. The optimal attack strategy for SHE is to predict the user’s value by selecting the index corresponding to the maximum component of the obfuscated vector: $\hat{x}=\underset{i\in[k]}{\mathrm{argmax}}\;y_{i}$ [31]. Following this attack strategy, we deduce the expected ASR for the SHE protocol as the probability that the noisy value $y_{x}$ at the true index exceeds all other noisy values $y_{i}$ for $i\neq x$:
$\text{ASR}_{\mathrm{SHE}}=\Pr\left[y_{x}>\max_{i\neq x}y_{i}\right].$ (21)
We defer the derivation of Eq. (21) to Appendix -A2.
IV-E2 Thresholding with HE (THE)
In THE, each perturbed component of the vector $\mathbf{y}$ is compared to a threshold value $\theta$ to generate the final output vector. More precisely:
$\forall{i\in[k]}:\quad\mathbf{y}_{i}=\begin{cases}1,&\text{if }\mathbf{y}_{i}>\theta\\ 0,&\text{if }\mathbf{y}_{i}\leq\theta\end{cases}$
Thus, the resulting output vector $\mathbf{y}$ is a binary vector in $\{0,1\}^{k}$, where we have the following probabilities:
$p$ $=\Pr[\mathbf{y}_{i}=1\mid\mathbf{x}_{i}=1]=1-\frac{1}{2}e^{\frac{\varepsilon}{2}(\theta-1)}\mathrm{.}$ (22)
$q$ $=\Pr[\mathbf{y}_{i}=1\mid\mathbf{x}_{i}=0]=\frac{1}{2}e^{-\frac{\varepsilon}{2}\theta}\mathrm{.}$ (23)
Aggregation. A reported bit vector $\mathbf{y}$ is viewed as supporting an input $i$ if $\mathbf{y}_{i}>\theta$. Therefore, the support set for THE is $\mathbb{1}_{\mathrm{THE}}=\{i\hskip 2.84544pt|\hskip 2.84544pt\textbf{y}_{i}\hskip 2.84544pt>\hskip 2.84544pt\theta\}$. The THE mechanism is pure with $p^{*}=p$ and $q^{*}=q$. The server estimates the frequency using Eq. (2) with the following analytical MSE:
$\mathrm{MSE}_{\mathrm{THE}}=\frac{2e^{\varepsilon\theta/2}-1}{n(1+e^{\varepsilon(\theta-1/2)}-2e^{\varepsilon\theta/2})^{2}}\mathrm{.}$ (24)
The optimal threshold value that minimizes the protocol’s MSE in Eq. (24) is within $\theta\in(0.5,1)$ [23].
Attacking. Based on the support set $\mathbb{1}_{\mathrm{THE}}$, the adversary can use one of two attack strategies denoted by ${\mathcal{A}}_{\mathrm{THE}}$ [31]:
  • ${\mathcal{A}}^{0}_{\mathrm{THE}}$ is a random choice $\hat{x}=\mathrm{Uniform}\left([k]\right)$, if $\mathbb{1}_{\mathrm{THE}}=\emptyset$;
  • ${\mathcal{A}}^{1}_{\mathrm{THE}}$ is a random choice $\hat{x}=\mathrm{Uniform}\left(\mathbb{1}_{\mathrm{THE}}\right)$, otherwise.
In this paper, we obtained the expected ASR for THE as:
$\mathrm{ASR}_{\mathrm{THE}}=(1-p)(1-q)^{k-1}\cdot\frac{1}{k}+\frac{p}{kq}\left(1-(1-q)^{k}\right)\mathrm{.}$ (25)
We defer the derivation of Eq. (25) to Appendix -A3.
V Attack-Aware Evaluation Framework for LDP
The existing literature has traditionally proposed mechanisms like SS, OUE, OLH and THE, with the primary objective of minimizing estimation error for a given privacy budget $\varepsilon$. However, protocols offering identical $\varepsilon$-LDP guarantees can exhibit substantially different vulnerabilities to inference and poisoning attacks, leading to markedly different real-world privacy and integrity risks. Motivated by this observation, we propose an attack-aware evaluation framework for LDP frequency estimation protocols, designed to systematically analyze how estimation utility interacts with adversarial vulnerability under different threat models and protocol configurations.
V-A General Evaluation Framework
We evaluate LDP protocols through a unified attack-aware instantiation, in which each mechanism is characterized by its behavior along two complementary axes beyond $\varepsilon$, capturing key system desiderata:
$\left(\mathcal{A}(\theta),\;\mathcal{U}(\theta)\right),$ (26)
in which $\theta\in\Theta$ denotes the protocol parameters (e.g., $\varepsilon$ or design choices), $\mathcal{U}$ measures utility loss and $\mathcal{A}$ quantifies the protocol attackability under a specified adversarial model. Specifically, the attackability component $\mathcal{A}$ can be instantiated to capture different classes of threats:
  • Privacy attackability ($\mathcal{A}_{\mathrm{priv}}$), including data reconstruction [30, 31], attribute inference [29, 10, 35] or re-identification [28, 27];
  • Integrity attackability ($\mathcal{A}_{\mathrm{sec}}$), such as vulnerability to data poisoning [34] or adversarial manipulation [41].
Utility loss $\mathcal{U}$ is typically quantified using metrics based on estimation error, such as MSE, Mean Absolute Error (MAE) [20], or Fisher Information [42].
This formulation provides a principled lens for reasoning about the inherent trade-offs between estimation accuracy and adversarial vulnerability in the design of LDP protocols.
V-B Specialization to Data Reconstruction and MSE Trade-off
In this work, we focus on a widely applicable and analytically tractable instantiation that captures the fundamental privacy-utility tension in LDP. Specifically, we consider:
  • Privacy attackability, measured via the ASR under data reconstruction attacks in Eq. (8). Reconstruction attacks directly target users’ private values and admit closed-form analyses for all eight state-of-the-art LDP protocols considered in this paper. Moreover, successful reconstruction can serve as a proxy for other downstream privacy risks [36, 37], such as profiling and re-identification (see Example 1) as well as attribute inference (see Example 2), making ASR a natural and interpretable indicator of broader privacy leakage.
  • Utility, measured by the variance (MSE) of frequency estimation, which directly quantifies estimation accuracy and is central to frequency estimation tasks [32, 23, 33]. Minimizing MSE also implicitly controls other utility losses such as MAE and Fisher Information through established theoretical relationships like the Central Limit Theorem and Cramér-Rao bounds [43, 44, 45].
This ASR-MSE instantiation yields a meaningful privacy-utility frontier that is central to many practical LDP deployments, and leverages closed-form expressions available for a wide range of protocols (see Section IV).
Protocol Selection via Pareto Optimization. Since ASR and MSE represent inherently competing objectives, each LDP protocol can be characterized by a set of Pareto-efficient configurations that balance privacy leakage and utility. To identify these configurations, we evaluate both metrics across a predefined parameter grid [46], forming the analytical Pareto frontier that captures the achievable trade-offs for each mechanism. Once the Pareto frontier is obtained, we select representative operating points using established multi-objective selection criteria (e.g., Utopia Point, Weighted Scalarization, Elbow “Knee” Method, etc) [47]. Among these, weighted scalarization stands out as a practical default due to its conceptual simplicity, interpretability and tunable nature. Empirically, we observed that all selection methods, except for the Elbow and $\boldsymbol{\epsilon}$-Constraint methods, tend to identify similar parameters (see Figure 7 in Appendix -B). For clarity and consistency, we therefore report results using the scalarized formulation:
$\underset{\mathbf{\Theta}}{\mathrm{min}}$ $J(\mathbf{\Theta})=w_{\text{ASR}}\cdot\text{ASR}+w_{\mathrm{MSE}}\cdot\mathrm{MSE}\mathrm{,}$ (27)
s.t. $w_{\text{ASR}}+w_{\mathrm{MSE}}=1\mathrm{.}$
Varying the weights $(w_{\text{ASR}},w_{\mathrm{MSE}})$ enables controlled exploration of privacy-utility preferences: higher $w_{\text{ASR}}$ emphasizes stronger protection, while larger $w_{\mathrm{MSE}}$ favors utility. This parameterization provides a unified way to visualize trade-offs across protocols and privacy regimes. Importantly, setting $w_{\text{MSE}}=1$ recovers the original protocols, showing that our refined mechanisms generalize rather than replace their classical counterparts.
Remark 1 (Attack-Agnostic Nature of the Framework).
The proposed attack-aware framework in Section V-A is not restricted to reconstruction attacks. More precisely, it is attack-agnostic and can accommodate arbitrary adversarial models, provided that the corresponding attackability metric can be evaluated, either analytically or via empirical (Monte Carlo) estimation.
In Section VI-E, we illustrate this flexibility by incorporating integrity-oriented attacks through data poisoning attacks [34]. While different attack metrics may interact with protocol parameters in distinct ways, the framework naturally supports their joint analysis, either via complementary two-dimensional trade-offs or through higher-dimensional visualizations and constrained selection strategies.
V-C Refining LDP Protocols
Using our two-objective framework defined in Eq. (27), we extend four state-of-the-art LDP protocols to introduce adaptive counterparts. Each adaptive protocol selects an optimal parameter $\mathbf{\Theta}$ to achieve a better trade-off between ASR and MSE, rather than focusing solely on utility. The optimization process varies across protocols, adapting their internal parameters to balance privacy protection and estimation accuracy. Importantly, our reparametrization preserves the original $\varepsilon$-LDP guarantee by design. For instance, the subset size $\omega$ in ASS and the probability pair $(p,q)$ in AUE are explicitly constrained to satisfy the $\varepsilon$-LDP guarantee; ALH inherits the privacy guarantee of GRR applied to a hashed domain; and ATHE applies a post-processing threshold, which does not affect the privacy guarantee. We now describe each adaptive mechanism in detail.
V-C1 Adaptive Subset Selection (ASS)
The ASS mechanism extends the SS protocol. Unlike SS, which aims to minimize the estimation error alone when selecting $\omega$, ASS jointly optimizes it for both MSE and ASR. Formally, the optimization problem for ASS is defined as:
$\begin{aligned} &\underset{\omega\in\mathbb{Z}}{\mathrm{min}}&&J(\omega)=w_{\mathrm{MSE}}\cdot\Bigg(\frac{(k-1)(k+2\omega e^{\varepsilon}-\omega)}{n\omega\left(-k+\omega+(k-1)e^{\varepsilon}-(\omega-1)e^{\varepsilon}\right)^{2}}\\ &&&\left.+\frac{\left(k-\omega+(\omega-1)e^{\varepsilon}\right)\left(-\omega(k-\omega)-\omega(\omega-1)e^{\varepsilon}\right)}{n\omega\left(-k+\omega+(k-1)e^{\varepsilon}-(\omega-1)e^{\varepsilon}\right)^{2}}\right)\\ &&&+w_{\text{ASR}}\cdot\left(\frac{e^{\varepsilon}}{\omega e^{\varepsilon}+k-\omega}\right),\\ &\emph{s.t.}&&1\leq\omega<k,\quad w_{\text{ASR}}+w_{\mathrm{MSE}}=1\mathrm{.}\end{aligned}$
The range $1\leq\omega<k$ ensures that at least one value is selected while keeping the subset smaller than the total domain size. This prevents trivial cases where $\omega=k$ would result in full randomness in the response or $\omega=0$ (no report).
V-C2 Adaptive Unary Encoding (AUE)
The AUE mechanism is a generalization of UE protocols. Unlike the optimized UE protocol (i.e., OUE [23]), which only aims to minimize the estimation error setting a fixed $p=1/2$ and $q=\frac{1}{e^{\varepsilon}+1}$, AUE jointly optimizes both MSE and ASR by adapting the probabilities $p$ and $q$. Formally, the optimization problem for AUE is defined as:
$\underset{p\in\mathbb{R}}{\mathrm{min}}$ $J(p)=w_{\text{MSE}}\cdot\left(\frac{\left(\left(e^{\varepsilon}-1\right)q+1\right)^{2}}{n\left(e^{\varepsilon}-1\right)^{2}(1-q)q}\right)$ (29)
$+w_{\text{ASR}}\cdot\frac{1}{k}\left[(1-p)(1-q)^{k-1}+\frac{p}{q}\left(1-(1-q)^{k}\right)\right]\mathrm{,}$
[…]
The equality constraint for $q$ is due to the $\varepsilon$-LDP requirement for UE protocols: $\varepsilon=\ln\left(\frac{p(1-q)}{(1-p)q}\right)$ [3, 23].
V-C3 Adaptive LH (ALH)
The ALH mechanism extends the LH protocol. Unlike the optimized LH protocol (i.e., OLH [23]), which aims solely to minimize estimation error by selecting $g=\lfloor e^{\varepsilon}+1\rceil$, ALH jointly optimizes both MSE and ASR by adapting the hash domain size parameter $g$. Formally, the optimization problem for ALH is defined as:
$\underset{g\in\mathbb{Z}}{\mathrm{min}}$ $J(g)=w_{\text{MSE}}\cdot\left(\frac{(e^{\varepsilon}-1+g)^{2}}{n(e^{\varepsilon}-1)^{2}(g-1)}\right)$ (30)
$\quad+w_{\text{ASR}}\cdot\left(\frac{e^{\varepsilon}}{(e^{\varepsilon}+g-1)\cdot\max\left\{\frac{k}{g},1\right\}}\right)\mathrm{.}$
s.t. $2\leq g\leq\max\left(k,\lfloor e^{\varepsilon}+1\rceil\right),\;w_{\text{ASR}}+w_{\mathrm{MSE}}=1\mathrm{,}$
The upper bound for $g$ is set to $\max\left(k,\lfloor e^{\varepsilon}+1\rceil\right)$, ensuring that $g$ is not unnecessarily smaller than the original domain size $k$. This avoids under-hashing, allowing for better differentiation and randomness. When $k$ is large, $g$ should ideally match or exceed $k$, ensuring that hashing effectively introduces randomness to maintain privacy guarantees.
V-C4 Adaptive THE (ATHE)
The ATHE mechanism extends the THE protocol. Unlike THE [23], which only aims to minimize the MSE in Eq. (24), ATHE jointly optimizes both MSE and ASR by adapting the threshold parameter $\theta$. Formally, the optimization problem for ATHE is defined as:
$\underset{\theta\in\mathbb{R}}{\mathrm{min}}$ $J(\theta)=w_{\mathrm{MSE}}\cdot\left(\frac{2e^{\varepsilon\theta/2}-1}{n(1+e^{\varepsilon(\theta-1/2)}-2e^{\varepsilon\theta/2})^{2}}\right)$ (31)
$\quad+w_{\text{ASR}}\cdot\Bigg((1-p)(1-q)^{k-1}\cdot\frac{1}{k}+\frac{p}{kq}\left(1-(1-q)^{k}\right)\Bigg)\mathrm{,}$
[…]
in which $p$ and $q$ are given in Eq. (22). The constraint $0.5\leq\theta\leq 1$ follows the settings used in prior work [23].
V-D Empirical Validation of Analytical ASR-MSE Frontiers
To validate the analytical ASR-MSE framework introduced in Section V-B, we empirically verify that the closed-form ASR and MSE expressions produce Pareto frontiers consistent with the observed behavior on real data. Concretely, we use the Adult dataset from the UCI machine learning repository [48], selecting the Age attribute with domain size $k=100$ (i.e., $\text{Age}\in[0,99]$) and $n=48842$ users, and report results averaged over $100$ independent runs. This experiment serves as a consistency check of the analytical methodology.
[…]
Figure 1 presents the comparison between empirical ($\circ$ markers) and analytical (red dashed lines) Pareto frontiers for ASR versus MSE across various LDP protocols, including both state-of-the-art and our adaptive variants. Each subplot represents a specific protocol (i.e., GRR, SUE, BLH, SHE, SS, OUE, OLH, THE, ASS, AUE, ALH and ATHE) with the privacy budget varying as $\varepsilon\in(2,10)$ with the domain size fixed at $k=100$. One can observe that across all protocols, the empirical results closely align with the analytical calculations, demonstrating the robustness of the closed-form ASR and MSE expressions. This alignment highlights the reliability of our theoretical framework and previous analytical experiments across different privacy budgets and data distributions.
Analytical vs. empirical validation: Our findings demonstrate a strong agreement between analytical and empirical ASR-MSE Pareto frontiers, supporting the use of closed-form expressions as a reliable basis for both systematic analysis and attack-aware LDP protocol design.
VI Experiments and Analysis
The objective of our experiments is to thoroughly evaluate the performance of the proposed adaptive LDP protocols in terms of privacy, utility and resilience against adversarial threats. Specifically, in Section VI-B, we conduct an ASR analysis for each LDP protocol to quantify their vulnerability to data reconstruction attacks, thereby assessing the privacy guarantees offered by existing and newly proposed mechanisms. Subsequently, in Section VI-C, we perform an MSE analysis to evaluate the utility guarantees provided by existing and our newly proposed mechanisms. Next, in Section VI-D, we explore the trade-off between ASR and MSE (i.e., Pareto frontier), highlighting how our adaptive protocols compare to traditional ones in balancing privacy and utility under different scenarios. Finally, in Section VI-E, we extend our evaluation to a joint view of privacy leakage, estimation utility and integrity under data poisoning attacks.
To systematize these evaluations, we rely on analytical closed-form expressions, enabling a comprehensive exploration of the privacy-utility-attackability trade-off across privacy budgets and domain sizes at low computational cost. This choice is supported by prior work showing close agreement between analytical and empirical results [23, 30, 13], and further validated by our empirical study in Section V-D.
VI-A Setup of Analytical Experiments
For all experiments, we have used the following settings:
  • LDP protocols. We experiment with the eight LDP protocols described in Section IV and our four adaptive LDP protocols described in Section V-C.
  • Number of users. For the analytical variance/MSE derived from Eq. (4) (e.g., Eq. (10) for GRR, Eq. (24) for THE, …), we report variance per user, corresponding to the analytical derivation of $\text{Var}[]/n$. This allows us to examine the fundamental properties of each protocol independently of the dataset size.
  • Privacy parameter. The LDP frequency estimation protocols were evaluated under two privacy regimes:
    1. a)
      High privacy regime with $\varepsilon\in\{0.1,0.2,\ldots,1.9,2.0\}$.
    2. b)
      Medium to low privacy regime with $\varepsilon\in\{2.0,2.5,\ldots,9.5,10.0\}$.
  • Domain size. We vary the domain size in three ranges:
    1. a)
      Small domain: $k\in\{25,50,75,100\}$.
    2. b)
      Medium domain: $k\in\{250,500,750,1000\}$.
    3. c)
      Large domain: $k\in\{2500,5000,7500,10000\}$.
  • Weights for optimization. Unless otherwise mentioned, we fix the weights for the two-objective optimization in Eq. (27) to $w_{\text{ASR}}=w_{\mathrm{MSE}}=0.5$, aiming for a balanced trade-off between privacy and accuracy. Experiments in Appendix -C focus on varying these weights.
  • Optimization method. Parameters for our adaptive protocols (e.g., subset size $\omega$ for ASS) were tuned via grid search [46], followed by selection using a Weighted Scalarization strategy (see Eq. (27)). While grid search is exhaustive and ensures broad coverage of the parameter space, it can be computationally expensive. For large or high-dimensional search spaces, we recommend constrained or gradient-free optimization methods [49] as a more efficient alternative to accelerate convergence without significantly sacrificing solution quality. Experiments in Appendix -D analyze parameter selection.
VI-B ASR Analysis for LDP Protocols
In Figure 2, we evaluate the ASR for various LDP frequency estimation protocols as a function of the privacy budget $\varepsilon$ from high to low privacy regimes across small to big domain sizes $k$. The goal of this analysis is to assess the robustness of each protocol against adversarial inference attacks under varying privacy levels and domain sizes. Specifically, we compare state-of-the-art protocols (GRR, SUE, BLH, OUE, OLH, SS, SHE and THE) against our proposed adaptive protocols (ASS, AUE, ALH and ATHE) to understand how effective these methods are at balancing privacy and utility.
Figure 2: Attacker Success Rate (ASR) vs. privacy budget ($\varepsilon$) for different LDP frequency estimation protocols across varying domain sizes ($k$). The plots compare state-of-the-art LDP protocols, including GRR, SUE, BLH, OUE, OLH, SS, SHE and THE, against our newly proposed adaptive protocols (ASS, AUE, ALH and ATHE). Each curve represents a different domain size, with $k$ ranging from $25$ to $10000$.
For all protocols, we observe that the ASR generally increases with increasing privacy budget $\varepsilon$ in Figure 2. This behavior is expected, as higher values of $\varepsilon$ correspond to weaker privacy guarantees, allowing the adversary to infer user data more effectively. For smaller domain sizes (i.e., $k\leq 100$), the ASR rises sharply, suggesting that the adversary’s ability to correctly guess the user’s input improves significantly as $\varepsilon$ increases. In contrast, larger domain sizes (i.e., $k\geq 1000$) show a more gradual increase in ASR, indicating that a larger domain inherently offers greater privacy protection. Still, our adaptive methods effectively counteract privacy threats regardless of $k$ and $\varepsilon$, underscoring the flexibility and robustness offered by our two-objective optimization framework.
Our adaptive protocols (i.e., ASS, AUE, ALH and ATHE) demonstrate significantly lower ASR compared to their traditional counterparts across all privacy budgets. For example, ASS effectively mitigates the ASR vulnerability of the state-of-the-art SS by capping the $\mathrm{ASR}$ below 0.25, in contrast to the traditional SS, where the $\mathrm{ASR}$ gets to $1$ (i.e., the adversary can fully infer the user’s value). Similarly, AUE consistently achieves a lower or comparable ASR relative to the state-of-the-art OUE protocol, highlighting the benefits of its adaptive parameter optimization. For ALH, the adaptive mechanism achieves a balanced compromise between the low ASR of BLH and the moderate ASR of OLH by dynamically optimizing the hash domain size, resulting in a substantial reduction in ASR compared to traditional OLH. Finally, ATHE demonstrates more resilience than THE across all domain sizes, thereby showcasing the effectiveness of our framework in enhancing privacy protection.
ASR increases with $\varepsilon$, but our adaptive protocols resist: As the privacy budget $\varepsilon$ increases, ASR generally rises for all protocols due to weaker privacy guarantees. However, our adaptive protocols (ASS, AUE, ALH and ATHE) exhibit significantly lower (i.e., by up to $\leq 5\times$) ASRs across a broad range of $\varepsilon$ values, underscoring their enhanced resilience against privacy attacks.
VI-C MSE Analysis for LDP Protocols
In Figure 3, we examine the MSE behavior of the SS protocol and our adaptive protocol ASS, across varying privacy budgets $\varepsilon$. In Figure 4, we extend this analysis to compare MSE trends for UE-, LH- and HE-based LDP protocols, including our adaptive versions (AUE, ALH and ATHE). These analyses aim to determine whether the adaptive protocols’ improved robustness to privacy attacks results in substantial estimation accuracy loss or if they maintain competitive MSE values. Notice that the MSE curves of SUE, OUE, BLH, OLH, SHE and THE remain independent of the domain size $k$ due to their fixed parameterization, as established in [23].
Figure 3: Variance (MSE) vs. privacy budget ($\varepsilon$) for the state-of-the-art SS protocol and our adaptive version ASS across various domain sizes $k$. Each curve represents a distinct domain size, illustrating how each protocol balances estimation accuracy with privacy as $\varepsilon$ changes.
Figure 4: Variance (MSE) vs. privacy budget ($\varepsilon$) for the state-of-the-art LDP protocols (UE-, LH-, and HE-based) and our adaptive versions (AUE, ALH, and ATHE) across various domain sizes $k$. For our adaptive protocols, each curve represents a distinct domain size, illustrating how each protocol balances estimation accuracy with privacy as $\varepsilon$ changes.
We observe in Figure 3 that ASS exhibits an increase in MSE by up to two orders of magnitude compared to SS, particularly in higher privacy regimes ($\varepsilon\geq 4$). This increase is expected, as SS’s minimal MSE comes at the cost of extreme vulnerability, with ASR approaching 1 (e.g., see Figure 2), meaning an adversary can fully infer the user’s value. ASS, in contrast, balances this trade-off by introducing adaptive parameterization that mitigates privacy attacks while moderately increasing the MSE. In Figure 4, we observe similar trends for the other adaptive protocols (AUE, ALH, and ATHE). For UE protocols, our adaptive AUE version achieves slightly higher MSE compared to OUE across all domain sizes $k$, with the gap becoming more pronounced as $\varepsilon$ increases. This is expected, as AUE optimizes its parameters to enhance robustness to privacy attacks, leading to a slight trade-off in utility. Notably, AUE remains competitive with SUE in high privacy regimes ($\varepsilon\leq 2$), offering similar MSE levels while achieving significantly improved ASR performance as shown earlier. In low privacy regimes ($\varepsilon>2$), AUE incurs a modest increase in MSE (1 order of magnitude) compared to OUE.
Moreover, for LH- and HE-based protocols, we observe in Figure 4 that our adaptive protocols (ALH and ATHE) demonstrate a variance (MSE) behavior that is consistently “sandwiched” between the two corresponding state-of-the-art protocols in their respective groups. Specifically, ALH achieves MSE values between BLH (which minimizes ASR at the cost of higher MSE) and OLH (which minimizes MSE but is more vulnerable to privacy attacks). Similarly, ATHE’s variance lies between SHE and THE, showing a trade-off where ATHE retains competitive MSE while prioritizing adversarial resilience. This positioning highlights how adaptivity allows our protocols to achieve a better privacy-utility trade-off.
Adaptive protocols maintain competitive MSE: Our adaptive protocols (ASS, AUE, ALH and ATHE) achieve higher MSE compared to their non-adaptive counterparts. However, these increases remain within acceptable bounds ($\leq 2$ orders of magnitude), highlighting that the improved adversarial resilience (ASR $\leq 5\times$) comes at a reasonable cost to the utility. This suggests that enhanced privacy protection against adversaries can be achieved without prohibitive increases in variance.
VI-D Pareto Frontier for ASR and MSE
Beyond analyzing ASR and MSE independently as functions of the privacy budget $\varepsilon$ (see Figures 24 and 3), it is crucial to examine how privacy leakage and estimation accuracy interact jointly. To this end, we study the Pareto frontier between attackability (ASR) and estimation error (MSE). Figure 5 reports ASR vs. MSE for a fixed domain size $k=100$, with the privacy budget $\varepsilon$ varying in the medium-to-low privacy regime. Protocols are grouped by their encoding family (GRR/SS, UE, LH and HE), allowing for direct comparison between state-of-the-art mechanisms and their adaptive counterparts under identical domain and privacy conditions. Each point corresponds to a specific value of $\varepsilon$, with colors indicating the privacy level. Additional per-protocol ASR-MSE trade-offs for other domain sizes and privacy regimes are provided in Appendix -E.
Figure 5: Attacker Success Rate (ASR) vs. Variance (MSE) for LDP frequency estimation protocols under data reconstruction attacks, with fixed domain size $k=100$. Each point corresponds to a distinct privacy budget $\varepsilon$ (medium to low privacy regimes), with colors indicating the privacy level. Protocols are organized by encoding family, comparing state-of-the-art mechanisms (GRR, SUE, BLH, SHE, SS, OUE, OLH, THE) against our adaptive counterparts (ASS, AUE, ALH, ATHE).
From Figure 5, comparisons within each encoding family reveal systematic structural differences in how protocols navigate the ASR-MSE trade-off. GRR and SS exhibit steep trajectories: as $\varepsilon$ increases, they rapidly reach low-MSE regimes at the cost of sharply increasing ASR, indicating strong utility but severe privacy leakage. In contrast, SUE-, SHE- and THE-based mechanisms follow smoother ASR-MSE paths, yielding more gradual trade-offs and avoiding abrupt transitions into highly vulnerable regions. OUE and OLH achieve low variance primarily at higher $\varepsilon$ values, where ASR plateaus around $0.5$, reflecting their utility-oriented parameterization. By contrast, BLH consistently maintains low ASR across the entire range of $\varepsilon$, but does so at the expense of substantially higher MSE, illustrating a conservative privacy-first design. Notably, because OUE is designed to minimize MSE, SUE achieves a more favorable ASR-MSE trade-off across a broad range of privacy budgets. This illustrates that minimizing estimation error alone does not necessarily yield the best privacy-utility balance under reconstruction attacks.
Our adaptive protocols fundamentally reshape these within-family trade-offs. ASS mitigates the sharp ASR escalation observed in SS at low-MSE operating points. AUE improves upon OUE by maintaining a lower ASR without incurring disproportionate variance increases, particularly in moderate privacy regimes, and in several regimes, it overlaps with SUE in the ASR-MSE space. Likewise, ALH and ATHE consistently dominate their non-adaptive counterparts by jointly reducing ASR without incurring excessive variance increases. Rather than interpolating between existing designs, these adaptive mechanisms shift the ASR-MSE trade-off frontier itself, enabling operating points that are unattainable through varying $\varepsilon$ alone under fixed protocol parameterizations.
Key insight: Each protocol exhibits a distinct ASR-MSE profile. By jointly optimizing for both ASR and MSE, our adaptive protocols consistently push these curves toward more favorable regimes in the ASR-MSE Pareto frontier.
VI-E Complementary Adversarial Perspective: Poisoning Attacks
While ASR captures vulnerability to data reconstruction attacks, LDP protocols may also be susceptible to data poisoning attacks [34], in which an adversary injects malicious reports to bias the aggregated frequency estimates. To capture this integrity-oriented threat (see Eq. (26)), we adopt the maximal gain attack (MGA) proposed by Cao et al. [34], which is shown to be the strongest poisoning strategy in terms of maximizing the expected increase in the estimated total frequency of a target item set $T$. For UE and LH protocols, the MGA admits a unified closed-form expression, referred to as the attacker’s Expected Poisoning Gain (EPG):
$\mathrm{EPG}=\beta\!\left(\frac{r(1-q)}{p-q}-f_{T}\right),$ (32)
in which $\beta$ denotes the fraction of fake users, $r=|T|$ is the number of targeted items and $f_{T}$ is the total true frequency mass of these items among genuine users. Following the experimental setup of [34], we fix the poisoning parameters to $\beta=0.05$, $r=1$, and $f_{T}=0.01$, ensuring comparability and avoiding additional attacker-side tuning degrees of freedom.
Joint View. Taken together, ASR, EPG and MSE define a three-dimensional trade-off between privacy leakage, estimation utility and integrity robustness. Figure 6 provides a geometric illustration of this joint space for state-of-the-art UE- and LH-based protocols and our adaptive counterparts, where each point corresponds to a distinct $\varepsilon\in(2,10)$ and a fixed $k=100$. A key implication is that decreasing $\varepsilon$ simultaneously increases the estimation error (MSE) and poisoning vulnerability (EPG), while reducing the reconstruction risk (ASR). Rather than prescribing a single operating point, this three-objective view supports multi-objective optimization or projections onto lower-dimensional trade-offs [47] (e.g., ASR vs. MSE). In our experiments, adaptive protocols are tuned exclusively with respect to ASR and MSE (Section V-C), while EPG is evaluated post hoc to assess the resulting integrity implications. More generally, alternative adaptive designs could directly incorporate poisoning robustness into the optimization objective, e.g., through weighted combinations of ASR, MSE and EPG as Eq. (27), which we leave as future work.
(a) UE-based protocols: SUE, OUE, and AUE.
(b) LH-based protocols: BLH, OLH, and ALH.
Figure 6: Joint ASR-MSE-EPG trade-off for UE- and LH-based LDP frequency estimation protocols at domain size $k=100$. Each point corresponds to a distinct privacy budget $\varepsilon\in(2,10)$.
Within this three-dimensional space, our adaptive protocols (AUE and ALH) reshape the attainable trade-offs at fixed $\varepsilon$. For UE-based protocols, AUE systematically bridges the gap between the utility-oriented OUE and the more conservative SUE, achieving comparable reconstruction rates to SUE at substantially lower estimation error. For LH-based protocols, ALH reduces ASR relative to OLH and, as a consequence of its parameter adaptation, also yields lower EPG, while achieving robustness levels comparable to BLH with substantially improved utility. Overall, these results demonstrate that attack-aware parameter adaptation reveals operating points that cannot be achieved by tuning the privacy budget alone.
VII Conclusion and Perspectives
This work presented a comprehensive experimental and analytical benchmark of LDP frequency estimation protocols, emphasizing the joint evaluation of privacy leakage, utility, and integrity under adversarial settings. By positioning existing protocols within a unified privacy-utility-attackability trade-off space, we provided a reproducible methodology to assess and compare their resilience against both data reconstruction and poisoning attacks. Our study revealed clear and protocol-dependent Pareto frontiers between ASR under data reconstruction attacks [30, 28, 31] and MSE, as well as their interaction with integrity-oriented attackability, exposing vulnerabilities not captured by $\varepsilon$ alone.
Beyond analysis, we showed that systematic benchmarking can guide the refinement of existing mechanisms. By leveraging multi-objective optimization, we derived adaptive parameterizations (ASS, AUE, ALH, ATHE) that achieve more favorable privacy-utility-attackability trade-offs than their original counterparts (SS, OUE, OLH, THE) while strictly preserving $\varepsilon$-LDP guarantees. These adaptive variants demonstrate that informed tuning can substantially enhance robustness against adversarial threats without prohibitive utility loss.
Overall, our work establishes a principled approach to evaluating LDP mechanisms through standardized, attack-aware benchmarks that jointly capture privacy, utility and integrity considerations. We envision this benchmark as a foundation for a broader suite of experimental tools for LDP, enabling future studies to extend our evaluation to additional adversarial models (e.g., attribute inference [29, 10, 35], re-identification [27, 28]) and complementary system-level metrics such as communication and computation cost.
Acknowledgments
The authors thank the anonymous reviewers for their insightful suggestions and Haoying Zhang for her helpful comments on an earlier draft. Héber H. Arcolezi has been partially supported by the French National Research Agency (ANR), under contracts: “ANR-24-CE23-6239” and “ANR-23-IACL-0006”. Sébastien Gambs is supported by the Canada Research Chair program as well as a Discovery Grant from NSERC.
AI-Generated Content Acknowledgement
The authors acknowledge the use of ChatGPT (OpenAI, GPT-5 model) to assist with language-related improvements, including grammar correction, spelling, formatting consistency and refinement of phrasing for clarity and readability. No part of the paper’s scientific content, analysis, experimental design, results or conclusions was generated or influenced by AI tools. All conceptual, technical and experimental contributions originate solely from the authors.